OpenStudy (anonymous):

What's 10 divided by 2/5 ?

OpenStudy (abb0t):

What's: \(\sf \color{red}{\frac{10}{1} \times \frac{5}{2}}\)?

OpenStudy (math_genius123):

put 2/5 as a decimal first

OpenStudy (math_genius123):

or do it his way

OpenStudy (math_genius123):

@abb0t way better and easier

OpenStudy (math_genius123):

25 bro 25